OpenStudy (anonymous):

Can someone help me with simplifying radicals? I'll comment my question I'm stuck on.

OpenStudy (anonymous):

This is what I've gotten so far: \[16\sqrt{343}=16\sqrt{49\times7}=16\sqrt{49}\times \sqrt{7}=272\sqrt{7}\] I am not sure if it's right! If I'm wrong, can you go through it with me?

OpenStudy (unklerhaukus):

You've started off right, factoring the term under the sqrt and breaking it up, but i think you made a multiplication error, [16×7≠272] \[16\sqrt{343}=16\sqrt{49\times7}=16\sqrt{49}\times \sqrt{7}=16\times7\times\sqrt7=...\sqrt{7}\]

OpenStudy (unklerhaukus):

just check that multiplication step 16×7 = ...

OpenStudy (anonymous):

Oh, I got that. So would the final answer be: \[112\sqrt{7}\]?

OpenStudy (unklerhaukus):

Yes!

OpenStudy (anonymous):

Okay, thank you for catching my mistake! I thought I got all of the steps right, but I had a feeling I made a mistake! @UnkleRhaukus

OpenStudy (unklerhaukus):

mistakes are easy to make , i still make (simple) multiplication errors all the time

OpenStudy (unklerhaukus):

one method to check this sort of problem is with a calculator [16][×][√(343)] = 296.3241468392341 [112]][×][√7] = 296.3241468392341

OpenStudy (anonymous):

Cool! That's nice to know!